OpenStudy (anonymous):

MEDAL+FAN If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ? (3 points) 12x + 9 12x + 7 12x + 5 12x + 3

OpenStudy (solomonzelman):

f(x) = 4x + 1 g(x) = 3x + 2 f(g(x))=4(3x+2) + 1 =12x+8 + 1 =12x+9

OpenStudy (anonymous):

Thank You

OpenStudy (solomonzelman):

You welcome !